I have a question. I am working on consider the following piecewise defined function. Evaluate the given at x = -5. Problem is f(x)= top part is 1/2x+6 if x is less than or equal to -5. Bottom one is -1/4x+7 if x is greater than -5 hope this makes sense!
Answer by Alan3354(69443)

Evaluate the given at x = -5. Problem is f(x)= top part is 1/2x+6 if x is less than or equal to -5. Bottom one is -1/4x+7 if x is greater than -5
------------------
Bottom one is -1/4x+7 if x > -5
That doesn't apply at x = -5, only > -5.
----------------
1/2x+6 if x <= to -5
(1/2)*-5 + 6 = 7/2